OpenStudy (anonymous):

A city park is opening a new swimming pool. You can pay a daily entrance fee of $3 or purchase a membership for the 12 week summer season for $82 dollars and pay only $1 per day to swim. How many days would you have to swim to make the membership worthwhile?

OpenStudy (anonymous):

did you get an answer to the last one?

OpenStudy (anonymous):

i got \(.0625\times 5280=330\)

OpenStudy (anonymous):

Yeah that's what I got

OpenStudy (anonymous):

k good this one is much easier

OpenStudy (anonymous):

for this one, put \(x\) as the number of days you go you will either pay \(3x\) dollars or \(82+x\) dollars solve \[3x=82+x\]

OpenStudy (anonymous):

pretty much do this one in your head right?

OpenStudy (anonymous):

subtract \(x\) get \[2x=82\] then take half of \(82\)

OpenStudy (anonymous):

So, x= 41 correct?